# Time-Scale Coupling Between States and Parameters in Recurrent Neural Networks

### v1

@@ 7 Conclusions and future directions
In this work, we have introduced a theoretical framework that couples the time scales of state-space dynamics with the parameter-space dynamics of recurrent neural networks. By modeling gating mechanisms as parametrized modifiers of state transitions, we demonstrated that their influence extends beyond state evolution: gates implicitly induce adaptive learning-rate behavior in parameter updates, even when training is carried out with a fixed global learning rate. Through a first-order Fréchet expansion of Jacobian products, we established how these gate-induced effects alter gradient propagation, scale step sizes, and introduce anisotropy in parameter updates.
Our derivations show that the corrections induced by gates can be treated as small, structured perturbations of the dominant dynamics in many practical regimes. Numerical experiments with constant, scalar, and multi-dimensional gates validated this approximation, confirming both the smallness of the corrections and the accuracy of the predicted scaling laws, even at the operating point $\varepsilon=1$.
Taken together, these results provide a unified dynamical-systems perspective on gating: gates serve not only as state filters but also as intrinsic mechanisms for adaptive optimization, effectively bridging state-space time scales and parameter-space learning dynamics. This perspective also highlights that adaptive learning-rate methods such as Adam or RMSprop could be revisited in light of these findings: rather than treating optimization and gating as separate mechanisms, future adaptive methods might explicitly incorporate the time-scale interactions revealed here, potentially leading to simpler or more principled training strategies.
Future work will extend this analysis to more advanced gated architectures such as LSTMs and GRUs, as well as modern sequence models with multiplicative interactions. We also plan to investigate the role of these time-scale interactions in tasks with long-range temporal dependencies, and to explore how the framework can guide the principled design of novel architectures that combine stability, training efficiency, and interpretability.
@@ Appendix A Matrix product expansion via the Fréchet derivative formulation
We derive here the first-order expansion of a product of matrices with structured perturbations, starting from the product rule for the Fréchet derivative (Theorem 3.3 in Higham [11]). We also show how the same reasoning yields higher-order terms.
@@ A.1 Fréchet Differentiability and the First-Order Expansion
Let $\mathbb{C}^{n\times n}$ denote the finite-dimensional vector space of complex $n\times n$ matrices equipped with a matrix norm (specifically, the operator 2-norm, unless otherwise stated). Since the space is finite-dimensional, all norms are equivalent, but fixing a concrete norm allows us to quantify perturbations in a consistent way.
Definition A.1 (Fréchet differentiability [11, Def. 3.1]).
@@
Intuitively, $L_{f}(A,\cdot)$ provides the best linear approximation to $f$ near $A$; for functions between finite-dimensional normed spaces, this is equivalent to the existence of the Gâteaux derivative that is uniformly continuous in a neighbourhood of $A$ [11, Sec. 3.1].
@@
where $E\in\mathbb{C}^{n\times n}$ is the perturbation matrix, specifying the direction and structure of the infinitesimal change applied to $A$. The notation $o(\|E\|)$ means that $\|o(\|E\|)\|/\|E\|\to 0$ as $\|E\|\to 0$.
@@
where $L_{g}(A,E)$ denotes the Fréchet derivative of $g$ at $A$ in the direction $E$; see [11, Sec. 3.2, Thm. 3.3] for a proof.
@@ A.4 First-order expansion
Higher-order terms
The same idea applies to higher derivatives: an $r$-th order term comes from choosing exactly $r$ distinct slots for the perturbations $B$. The general $r$-th Fréchet derivative at $\varepsilon=0$ is
$F^{(r)}(0)=r!\!\!\!\sum_{1\leq m_{1}<\cdots<m_{r}\leq n}\!\!\!\left(\prod_{j=m_{r}+1}^{n}A_{j}\right)\,B_{m_{r}}\left(\prod_{j=m_{r-1}+1}^{m_{r}-1}A_{j}\right)\,B_{m_{r-1}}\cdots B_{m_{1}}\left(\prod_{j=1}^{m_{1}-1}A_{j}\right).$ (76)
Thus the exact Taylor expansion (terminating at degree $n$) is
$F(\varepsilon)=\sum_{r=0}^{n}\frac{\varepsilon^{r}}{r!}F^{(r)}(0).$ (77)
Note that $F^{(1)}(0)\equiv L_{F}(0,E)$.

### v2

@@ 6 Analogies between gating mechanisms and adaptive gradient descent methods
7 Simulations
7.1 Effective learning rate induced by gates
We empirically verify that, even when SGD uses a fixed global step size $\mu$, gating mechanisms induce a lag-dependent effective learning rate that depends on the state-space model (leaky, scalar gate, multi-gate). The central object we use is the lag-conditioned sensitivity
$S_{t,k}(\theta)\;=\;\Big{\|}\,\prod_{j=k+1}^{t}J_{j}(\theta)\,\Big{\|}_{2},$
whose magnitude reweights how much a gradient signal at time $k$ influences the update due to a loss at time $t$. At each training checkpoint $\ell$ we aggregate these quantities to obtain an empirical effective learning rate profile. For every temporal distance h=tkh=t-k, we define
$\tilde{\mu}_{\mathrm{eff}}(h;\ell)\;\propto\;\mathrm{median}_{(t,k):\,t-k=h}\,S_{t,k}(\theta_{\ell}).$
To ensure comparability across models and checkpoints, we report a normalized version anchored at unit lag:
$\tilde{\mu}_{\mathrm{eff}}(h;\ell)\;=\;\frac{\mathrm{median}_{(t,k):\,h}\,S_{t,k}(\theta_{\ell})}{\mathrm{median}_{(t,k):\,h=1}\,S_{t,k}(\theta_{\ell})}\,.$
This construction emphasizes the relative attenuation with lag: all curves start from $1$ at $h=1$, so differences directly reflect how fast or slow the effective learning rate decays with temporal distance. The normalization removes global scale effects due to parameter growth and isolates the temporal structure of the effective learning rate.
Predictors and overlays.
To interpret the empirical effective learning rates, we compare them against a simple gate-product predictor derived from the zeroth-order theory of Section 5.2. For each model we define
$P_{t,k}(\theta_{\ell})\;=\;\begin{cases}\alpha^{\,t-k},&\text{leaky (constant $\alpha$)},\\[2.0pt] \prod_{j=k+1}^{t}g_{j-1},&\text{scalar gate},\\[6.0pt] \frac{1}{N_{r}}\displaystyle\sum_{i=1}^{N_{r}}\;\prod_{j=k+1}^{t}g^{(i)}_{j-1},&\text{multi-gate},\end{cases}$
which captures the multiplicative attenuation induced purely by the gates. We then form its lag-binned median
$\bar{P}(h;\ell)=\mathrm{median}_{(t,k):\,t-k=h}\,P_{t,k}(\theta_{\ell}).$
From this baseline predictor, we construct two reference curves to overlay on the empirical profiles:
$\tilde{\mu}^{(0)}_{\mathrm{pred}}(h;\ell)$ $=\frac{\bar{P}(h;\ell)}{\bar{P}(1;\ell)},$ (zeroth-order theory, no fitting)
$\tilde{\mu}^{(\mathrm{fit})}_{\mathrm{pred}}(h;\ell)$ $=\Bigg{(}\frac{\bar{P}(h;\ell)}{\bar{P}(1;\ell)}\Bigg{)}^{s(\ell)},$ $\text{(empirical correction)}.$
The first curve $\tilde{\mu}^{(0)}_{\mathrm{pred}}$ is the direct theoretical prediction obtained by normalizing the gate-product decay. The second curve $\tilde{\mu}^{(\mathrm{fit})}_{\mathrm{pred}}$ introduces a fitted exponent $s(\ell)$ that corrects for higher-order (perturbative) effects not captured by the pure gate product. The slope $s(\ell)$ is estimated at checkpoint $\ell$ via a robust log–log regression
$\log S_{t,k}\;\approx\;a(\ell)+s(\ell)\,\log P_{t,k},$
restricted to the central predictor range (1%–99% quantiles) to avoid extreme-tail artifacts.
Setup.
We train each model (leaky, scalar, multi-gate) with plain SGD (no momentum, no adaptivity) on a simple sequence task – adding problem. At selected checkpoints, we (i) run a fixed probe batch, (ii) build the exact one-step Jacobians $J_{j}$ from Section 5.1, (iii) estimate $S_{t,k}$ via a short power method iteration (5 steps) for the products $\prod_{j=k+1}^{t}J_{j}$, a standard technique to approximate the largest singular value, (iv) compute $P_{t,k}$, and (v) produce the normalized lag profiles with theory overlays. We also track the fitted slope $s(\ell)$ across training.
Results.
Across all models, the empirical profiles $\tilde{\mu}_{\mathrm{eff}}(h;\ell)$ exhibit a clear, monotone attenuation with lag and are well tracked by the fitted-power overlay $\tilde{\mu}^{(\mathrm{fit})}_{\mathrm{pred}}(h;\ell)$; the zeroth-order overlay $\tilde{\mu}^{(0)}_{\mathrm{pred}}(h;\ell)$ captures the right trend but is often too conservative or too aggressive depending on the model. The fitted log–log slopes extracted from the summaries (slope, $R^{2}$) are:
  • Leaky (constant $\alpha$): $s(\ell)$ starts around $2.17$ at iteration 0 and slowly decreases to $\approx 2.05$ by iteration $600$ (all $R^{2}\in[0.95,0.97]$). The effective profile attenuates more steeply than the pure $\alpha^{h}$ predictor, reflecting additional contraction from the perturbative contraction terms (identity leakage and activation derivatives).
  • Scalar gate: $s(\ell)$ is $\approx 0.30$ at iteration 0 and decreases toward $\approx 0.23$ by iteration $800$ ($R^{2}\approx 0.99$ throughout). The empirical profile decays more slowly than the product of gates, indicating that the $(1-g_{t})x_{t}$ pathway and gate-gradient terms partially counteract multiplicative attenuation, preserving longer-range contributions.
  • Multi-gate: $s(\ell)$ remains in a narrow band $\approx 0.45$$0.49$ from iteration 0 to $800$ ($R^{2}\approx 0.99$). The attenuation is sublinear relative to the gate product but stronger than in the scalar case. While the multi-gate architecture implies neuron-wise anisotropy in principle (each unit has its own gate trajectory), our scalar sensitivity measure $S_{t,k}$ only captures the dominant contraction rate. The fitted slope $s(\ell)$ should therefore be interpreted as an aggregate effect of this anisotropic gating.
Interpretation.
The slope $s(\ell)$ provides a quantitative diagnostic of how gates renormalize effective learning rates:
  • $s=0$: no lag dependence, i.e. the effective learning rate remains constant and equal to the nominal $\mu$.
  • $s=1$: exact agreement with the zeroth-order gate-product predictor, i.e. $\tilde{\mu}_{\mathrm{eff}}(h)\propto P_{t,k}$.
  • $s>1$: effective sensitivity decays more steeply with lag than the predictor, shortening the memory horizon and yielding more conservative updates.
  • $0<s<1$: effective sensitivity decays more slowly with lag than the predictor, extending the memory horizon and yielding more permissive updates.
The fact that $s(\ell)$ is neither 0 nor $1$ in practice confirms the presence of first-order correction terms beyond the pure gate product. Importantly, $s(\ell)$ remains relatively stable across training, suggesting that this modulation is a structural property of the dynamics rather than a transient effect of optimization.
Task dependence of the slope.
The fitted log–log slope $s(\ell)$ is not an architecture-only quantity: it depends on how the data and objective shape the Jacobians $J_{j}(\theta_{\ell})$ through the activation regime and the statistics of the gates. Tasks that keep units in a near-linear regime and encourage persistently open gates tend to yield $0<s<1$ (slower-than-predictor decay of the empirical effective learning rate with lag), whereas tasks that drive saturation or stronger filtering typically yield $s>1$ (faster-than-predictor decay). Practically, this means the naive gate-product predictor is renormalized by first-order effects in a task-specific manner. Conditioning the fit on gate-product or activation-slope quantiles leads to consistent shifts in $s$, and longer-horizon tasks generally reduce $s$ by encouraging memory-preserving dynamics. Thus, observed deviations of $s$ from $1$ quantify meaningful, task-induced corrections beyond the zeroth-order gate product and should be interpreted as part of the learned, problem-dependent time-scale interaction.
Figures.
We include, for each model, (left) a representative profile at the final checkpoint with both theory overlays, (middle) the evolution of the fitted slope $s(\ell)$ across training, and (right) the sensitivity heatmap $S_{t,k}$ showing the full temporal structure.
Figure 1: Leaky RNN (constant $\alpha$): normalized effective LR profile at final checkpoint (left), slope $s(\ell)$ across iterations (middle), and full sensitivity heatmap $S_{t,k}$ (right).
Figure 2: Scalar-gated RNN: normalized effective LR profile at final checkpoint (left), slope $s(\ell)$ across iterations (middle), and full sensitivity heatmap $S_{t,k}$ (right).
Figure 3: Multi-gated RNN: normalized effective LR profile at final checkpoint (left), slope $s(\ell)$ across iterations (middle), and full sensitivity heatmap $S_{t,k}$ (right).
7.2 Anisotropy of gradient propagation: gates vs. optimizer
In recurrent networks, the backpropagated error signal at time $k$ is obtained by transporting the loss gradient at time $t$ through the Jacobian products
$M_{t,k}\;=\;\prod_{j=k+1}^{t}J_{j}\,.$
These operators control not only the magnitude of gradients (how much signal survives with lag) but also their directional structure. Any gradient vector reaching earlier timesteps is linearly transformed by $M_{t,k}$; if the spectrum of $M_{t,k}$ is concentrated in a few large singular values, then error signals from many different losses will be funneled into the same dominant subspace. In this case, gradient flow becomes effectively low-dimensional, biasing parameter updates toward specific search directions. Conversely, if singular values remain comparable, gradients are spread across many directions, leading to more isotropic updates.
We therefore define anisotropy of gradient propagation as the degree of directional concentration induced by the singular spectrum
$\sigma_{1}(M_{t,k})\;\geq\;\cdots\;\geq\;\sigma_{N_{r}}(M_{t,k})\,.$
A rapid decay in $\sigma_{i}$ with $i$ indicates that long-range sensitivities align with a restricted set of directions, whereas a flatter spectrum reflects a more isotropic distribution of gradient energy across state space.
This section quantifies such anisotropy and compares three models: (i) a plain RNN trained with Adam (no gates, adaptive optimizer), (ii) a scalar-gated RNN trained with SGD (15), and (iii) a multi-gated RNN (16) trained with SGD. The central question is whether neuron-wise gates with a fixed global step size induce directional concentration of temporal sensitivities that is comparable to, or stronger than, optimizer-induced preconditioning.
Quantities.
For each probe sequence and checkpoint $\ell$, we form Jacobian products $M_{t,k}(\theta_{\ell})=\prod_{j=k+1}^{t}J_{j}(\theta_{\ell})$ and compute (via svd) its singular values $\sigma_{1}\!\geq\!\cdots\!\geq\!\sigma_{N_{r}}$. We aggregate over all pairs with lag h=(tk)h{=}(t{-}k) to obtain two lag-indexed metrics:
  • Anisotropy Index (AI): $\mathrm{AI}_{r}(h;\ell)=\mathrm{median}_{h}\big{(}\sigma_{1}/\sigma_{r}\big{)}$, which is large when sensitivity mass concentrates along a leading direction. We use r=8r{=}8.
  • Energy concentration (CE): $\mathrm{CE}_{r}(h;\ell)=\mathrm{median}_{h}\big{(}\sum_{i=1}^{r}\sigma_{i}^{2}/\sum_{i}\sigma_{i}^{2}\big{)}$, which lies in $[0,1]$ and is high if the top $r$ modes dominate. We use r=8r{=}8.
Both metrics summarize the directional geometry of temporal credit assignment: AI emphasizes the spread between the first and $r$-th mode; CE measures how much of the total sensitivity energy sits in the top subspace.
Setup.
We train the three models above on two standard sequence problems: (i) the adding task (real-valued, two marked positions), and (ii) a compact copy task (one-hot tokens followed by a delimiter, with a regression proxy target). All models have the same state dimension and are trained to comparable loss (plain RNN uses Adam, gated models use plain SGD; no momentum). At selected iterations $\ell\in\{0,400,800\}$ we run a fixed probe batch, build $J_{j}(\theta_{\ell})$, form products $M_{t,k}$ for lags $h\in\{1,2,4,8,12,16,24,32\}$, and record AI and CE per lag (median and interquartile range).
Results (adding task).
At initialization, the multi-gated model trained with SGD already exhibits pronounced directional concentration for medium and long lags, with a larger anisotropy index (AI $\sigma_{1}/\sigma_{r}$) than the plain RNN with Adam at several lags (Figure 4). As training proceeds, the plain RNN with Adam rapidly increases its anisotropy: by the final checkpoint (iter 800) it attains the highest AI across nearly all lags, whereas the multi-gated model remains consistently anisotropic but below Adam, and the scalar-gated model stays lowest among the three. The energy-concentration metric (CEr) mirrors these trends: CE rises with lag for all models, with Adam ultimately achieving the strongest concentration (largest CE) at the last checkpoint, multi-gate+SGD intermediate, and scalar gate weakest. Taken together, these curves indicate that optimizer-induced preconditioning (Adam) can produce stronger low-dimensional amplification of temporal sensitivities than gates alone on this task; nevertheless, neuron-wise gates with a fixed step size still induce substantial, structured anisotropy without any adaptive optimizer.
Results (copy task).
On the copy task, all three models display growing anisotropy with lag and training. The multi-gated RNN with SGD consistently achieves the highest anisotropy index across checkpoints and lags, indicating that per-neuron gating is particularly effective at biasing temporal dynamics toward dominant low-dimensional directions. The plain RNN trained with Adam does increase its anisotropy compared to initialization, but remains below the multi-gate baseline throughout training. Moreover, in the CEr plots Adam exhibits a distinct drop at larger lags (24 and 32), suggesting that the optimizer-induced concentration weakens when gradient propagation is stretched further in time. The scalar-gated model again provides only modest anisotropy, positioned between plain SGD and multi-gate but substantially below both at long horizons.
Takeaways.
(i) All models exhibit increasing anisotropy with lag, confirming that long-range gradient flow becomes effectively low-dimensional. (ii) Per-neuron gates with a fixed global step size yield the most robust and consistent anisotropy, outperforming Adam across training on the copy task. (iii) Adam can boost anisotropy relative to a plain RNN, but its effect is task-dependent: on the adding task it eventually surpasses the multi-gate baseline, whereas on the copy task it remains weaker and loses concentration at longer lags. (iv) The scalar gate provides some concentration but markedly less than multi-gate. These findings complement the results in Section 7.1: beyond scaling the magnitude of effective learning rates with lag, gates and optimizers also sculpt which directions in state space carry long-range gradient information, with multi-gate architectures proving especially effective for tasks requiring persistent memory.
Figure 4: Early training. Anisotropy is already detectable for multi-gate and grows with lag; the relative ordering across models is preserved throughout training.
Figure 5: Adding task. Multi-gate (SGD) exhibits the strongest directionality (higher AI, higher CE) across lags; scalar gate (SGD) is intermediate; plain RNN (Adam) is lowest. Error bands show interquartile ranges across $(t,k)$ pairs at each lag.
Figure 6: Copy task. Directional focusing with multi-gate is stronger and emerges at smaller lags than on the adding task; scalar gate remains intermediate; plain+Adam concentrates more gradually.
7.3 Broader implications
The simulations of Sections 7.1 and 7.2 reveal two complementary roles of gating and optimizer dynamics. First, gates modulate the magnitude of effective learning rates across lags by controlling how much past information is preserved in the state trajectory. Second, gates also shape the directional structure of gradient propagation, determining which subspaces of the state space remain active over long horizons.
These findings highlight that gating mechanisms do more than just stabilize training: they act as structural priors that bias the geometry of gradient flow. Optimizer adaptivity (e.g. Adam) can induce similar concentration effects, but the results show that this behavior is task-dependent: on the adding problem, Adam eventually matches or exceeds the anisotropy induced by multi-gating, while on the copy task its concentration weakens at longer lags. In contrast, per-neuron gating produces robust anisotropy across tasks. The scalar gate baseline sits between these extremes, offering partial control over anisotropy but without the full flexibility of neuron-wise gating.
Taken together, these results suggest a general design principle: architectures and optimizers jointly determine not only how fast parameters move, but also which temporal and spatial directions dominate learning. This dual role has direct relevance for modern sequence models. In gated RNNs, the anisotropy induced by gates may interact with optimizer preconditioning in subtle ways; in attention-based models, analogous phenomena may arise through value/attention weighting. Our framework provides a lens to study such interactions systematically, pointing toward hybrid strategies where structural gating and optimizer dynamics are co-designed to control both effective learning rates and anisotropy of temporal credit assignment.
@@ 8 Conclusions and future directions
We developed a theory of time-scale interaction between state and parameter dynamics in RNNs, showing how gating mechanisms not only control information flow in state-space, but also simultaneously regulate effective learning rates and the anisotropy of gradient propagation. Our simulations confirmed that: (i) per-neuron gates reliably extend the memory horizon and induce strong anisotropy, (ii) scalar gates offer intermediate control, (iii) adaptive optimizers such as Adam can mimic or even surpass gate-induced anisotropy in some tasks, but their behavior remains task-dependent, and (iv) the interplay between gates and optimizers determines both how much and which directions of past information shape parameter updates.
These findings suggest that gates should be understood also as structural priors for temporal geometry, while optimizers act as dynamic reweighting mechanisms. This perspective opens new avenues for co-designing architectures and optimizers to achieve desired trade-offs in memory, stability, and efficiency of sequence learning. Extensions to LSTMs [13] and GRUs [6] and transformers [34] constitute promising directions for future work.
@@ Appendix A Matrix product expansion via the Fréchet derivative formulation
We derive here the first-order expansion of a product of matrices with structured perturbations, starting from the product rule for the Fréchet derivative (Theorem 3.3 in Higham [12]).
@@ A.1 Fréchet differentiability and the first-order expansion
Let $\mathbb{C}^{n\times n}$ denote the finite-dimensional vector space of complex $n\times n$ matrices equipped with a matrix norm (specifically, the operator 2-norm, unless otherwise stated). Since the space is finite-dimensional, all norms are equivalent.
Definition A.1 (Fréchet differentiability [19, 12]).
@@
where $E\in\mathbb{C}^{n\times n}$ is the perturbation matrix, specifying the direction and structure of the infinitesimal change applied to $A$. The notation $o(\|E\|)$ means that $\|o(\|E\|)\|/\|E\|\to 0$ as $\|E\|\to 0$, i.e. the error must vanish at first order relative to the perturbation size.
@@
where $L_{g}(A,E)$ denotes the Fréchet derivative of $g$ at $A$ in the direction $E$; see [12, Sec. 3.2, Thm. 3.3] for a proof.
@@ A.5 Simulations supporting the validity of the first-order approximation
Appendix B Adam
$\hat{m}_{t,l}$ $=\frac{m_{t,l}}{1-\beta_{1}^{l}},$ (72)
$\hat{v}_{t,l}$ $=\frac{v_{t,l}}{1-\beta_{2}^{l}}.$ (73)
The parameter update is then: